# NFIG: Multi-Scale Autoregressive Image Generation via Frequency Ordering

### v3

@@ Abstract
Autoregressive models have achieved significant success in image generation. However, unlike the inherent hierarchical structure of image information in the spectral domain, standard autoregressive methods typically generate pixels sequentially in a fixed spatial order. To better leverage this spectral hierarchy, we introduce Next-Frequency Image Generation (NFIG). NFIG is a novel framework that decomposes the image generation process into multiple frequency-guided stages. NFIG aligns the generation process with the natural image structure. It does this by first generating low-frequency components, which efficiently capture global structure with significantly fewer tokens, and then progressively adding higher-frequency details. This frequency-aware paradigm offers substantial advantages: it not only improves the quality of generated images but crucially reduces inference cost by efficiently establishing global structure early on. Extensive experiments on the ImageNet-256 benchmark validate NFIG’s effectiveness, demonstrating superior performance (FID: 2.81) and a notable 1.25$\times$ speedup compared to the strong baseline VAR-d20.
@@ 3 Next-Frequency Image Generation
To provide a comprehensive understanding of our NFIG methodology, this section delves into its intricate architectural structure. The essential operational sequence, illustrating the flow and interaction of the system’s key components, is clearly visualized in Figure 2.
@@ 3.2 Autoregressive Image Generation
where $T_{i}\in[K]^{h_{i}\times w_{i}}$ is the matrix of code indices for the $i$-th frequency component, and the set $[K]=\{1,2,\dots,K\}$ represents all available index values.
@@ 4.1 Experimental Settings
Implementation Details. Our model was implemented using the PyTorch framework [42] and trained on NVIDIA H100 graphics cards. To ensure the reproducibility of our experiments, our implementation is built upon open-source research code, while incorporating improvements specific to this study. For the image tokenizer, the FR-VAE incorporates a VQGAN architecture with a DINO discriminator. The image encoder is initialized with pretrained weights from DINOv2-base. The frequency residual quantizer employs multiple scaling factors $[1,2,3,4,5,6,8,10,13,16]$ across different frequency bands, resulting in a vocabulary size of 680 tokens. The FR-VAE codebook size of 4096 was utilized. The image generator employs a VAR Transformer backbone with a depth of 16, enabling multi-scale image prediction. Optimization was performed using the Adam optimizer, setting the learning rate to $8\times 10^{-5}$ and the batch size to 768. Training of the model ran for 350 epochs on the ImageNet dataset. For inference, we configured CFG to 4.5 and top_k to 990.
@@ 4.2 Main Results
AR Model Comparison. NFIG achieves the best gFID (2.81) and IS (332.42) scores, significantly outperforming other AR models. Compared to VAR-16 (gFID: 3.55, IS: 274.4), our approach reduces FID by 0.74 and improves IS by more than 21%. Additionally, the proposed approach outperforms VAR-d20, a relatively larger model, while delivering 25% faster inference speed.
@@
Qualitative Results. Figure 3 qualitatively shows NFIG’s impressive ability to generate diverse ImageNet $256\times 256$ images across a wide variety of categories.
@@ 4.4 Motivation Verification
Frequency Distribution Analysis. The experimental results in Figure 4 demonstrate the progressive refinement of generated images and the effective capture and synthesis of multi-scale visual features by the NFIG model. The frequency spectrum visualizations reveal the model’s ability to hierarchically incorporate information from low to high frequencies, resulting in generated images with rich details and natural appearance.
Figure 4: Generated images at different steps 2, 4, 6, 8, 10 of a 10-step process, with corresponding frequency spectrum. In these spectrograms, brightness (red/yellow) indicates higher frequency energy while darker colors (blue) represent lower energy components. The center of each plot shows low-frequency information, with frequencies increasing radially outward, revealing the evolving distribution during the generation process.
@@ 6 Limitation and Future Work
Our frequency-guided autoregressive image generation approach shows promise, but has limitations. A primary issue is the simplistic frequency band division by scale, which inadequately captures information in the first band. Implementing a more rigorous division based on statistical analysis and physical principles would likely enhance NFIG’s performance. Beyond 2D spatial frequency, future work could extend to video generation, a more comprehensive direction that requires incorporating both temporal and spatial frequencies. Similarly, 3D object generation offers a better approximation of real-world light fields than 2D projections, where frequency analysis is vital for accurate representation. Due to computational constraints and time limitations, these promising directions remain unexplored and are left for future investigation.
@@ A.3 Magnitude Spectrum and Power Spectrum
Appendix B: Conceptual Parallel: Quantum Field Theory and Frequency-Based Image Generation
A Conceptual Parallel: Quantum Field Theory and Frequency-Based Image Generation – An Inspiring Perspective
Introduction
Quantum Field Theory (QFT) provides a fundamental description of the universe, viewing physical reality in terms of quantum fields rather than discrete point particles. Interestingly, this framework, particularly its treatment of fields and their excitations, presents some insightful conceptual parallels with modern frequency-based approaches to image generation, such as NFIG. Exploring this connection might offer new perspectives for both fields.
Core Concepts and Parallels
At its core, QFT sees the universe as composed of continuous quantum fields that fill all of spacetime. Particles, like electrons or photons, emerge as quantized excitations of these underlying fields. A key aspect is that these fields can be described as a collection of fundamental "modes," each characterized by specific energy and momentum. These quantities are linked to frequency and wavevector through the fundamental Planck-Einstein and de Broglie relations:
$\text{Energy}\quad E$ $=\hbar\omega\quad(\text{related to temporal frequency})$
@@ Introduction
where $\hbar$ is the reduced Planck constant, $\omega$ is angular frequency, and $\mathbf{k}$ is the wavevector. Thus, exciting a specific mode of the quantum field corresponds to creating a particle with a definite frequency and spatial variation.
Correspondingly, frequency-based image generation methods view an image as a composition of different spatial frequency components. Techniques like NFIG often build images hierarchically, starting with lower-frequency content to establish the coarse structure and overall composition, and then progressively adding higher-frequency components to introduce fine details.
Drawing on this, we can identify several conceptual parallels. Both QFT and frequency-based image generation involve the idea of decomposing a complex system (physical reality or an image) into fundamental constituents characterized by frequency or momentum/wavevector. In QFT, physical states are understood as arising from exciting field modes, while in image generation, images are built from spatial frequency components. Furthermore, both often involve a hierarchical structure: QFT relates different energy/momentum scales to distinct physical phenomena and interactions, while frequency bands in image generation capture image features at different spatial scales, from coarse structures (low frequencies) to fine details (high frequencies). Finally, just as complex quantum states are formed by superposing mode excitations, a complete image is formed by superimposing its various frequency components.
Potential Research Insights
Nevertheless, exploring this conceptual link might stimulate thinking and offer potential avenues for future research. One might consider whether mathematical structures or principles from QFT could inform the design of new image generation algorithms, particularly for multi-scale or frequency-space operations. Similarly, QFT’s approach to relating different energy scales to physical phenomena could potentially provide inspiration for learning more effective multi-scale representations within neural networks. More broadly, this perspective hints at the possibility of developing more unified theoretical frameworks that could potentially bridge concepts in fundamental physics and complex learned systems like neural networks.
Conclusion
Ultimately, this conceptual parallel serves as an inspiring thought framework rather than a strict equivalence. It highlights shared principles in how complex systems can be understood and constructed through their fundamental frequency/momentum components, potentially fostering interdisciplinary dialogue and novel approaches.

### v4

@@ Abstract
[…]
@@ 3 Next-Frequency Image Generation
To provide a comprehensive understanding of our NFIG methodology, this section delves into its intricate architectural structure. The essential operational sequence, illustrating the flow and interaction of the system’s key components, is clearly visualized in Figure 2. The details of loss function have been listed in Appendix B.1.
@@ 3.2 Autoregressive Image Generation
where $T_{i}\in[K]^{h_{i}\times w_{i}}$ is the matrix of code indices for the $i$-th frequency component, and the set $K]=\{1,2,\dots,K\}$ represents all available index values.
@@ 4.1 Experimental Settings
Implementation Details. Our model was implemented using the PyTorch framework [42] and trained on NVIDIA H100 graphics cards. To ensure the reproducibility of our experiments, our implementation is built upon open-source research code, while incorporating improvements specific to this study. For the image tokenizer, the FR-VAE incorporates a VQGAN architecture with a DINO discriminator. The image encoder is initialized with pretrained weights from DINOv2-base. Since VAR’s image tokenizer training code is not open-source, we adopted XQGAN’s implementation strategy. The frequency residual quantizer employs multiple scaling factors $[1,2,3,4,5,6,8,10,13,16]$ across different frequency bands, resulting in a vocabulary size of 680 tokens. The FR-VAE codebook size of 4096 was utilized. The image generator employs a VAR Transformer backbone with a depth of 16, enabling multi-scale image prediction. Optimization was performed using the Adam optimizer, setting the learning rate to $8\times 10^{-5}$ and the batch size to 768. Training of the model ran for 350 epochs on the ImageNet dataset. For inference, we configured CFG to 4.5 and top_k to 990.
@@ 4.2 Main Results
AR Model Comparison. NFIG achieves the best gFID (2.81) and IS (332.42) scores, significantly outperforming other AR models. Compared to VAR-16 (gFID: 3.55, IS: 274.4), our approach reduces FID by 0.74 and improves IS by more than 21%. XQGAN-310M has a better image tokenizer with rFID 0.78 with gFID 2.96. This indicates that NFIG’s performance improvement is not solely due to a good image tokenizer, but more importantly, to the injection of image frequency priors. Additionally, the proposed approach outperforms VAR-d20, a relatively larger model, while delivering 25% faster inference speed.
@@
Qualitative Results. Figure 3 qualitatively shows NFIG’s impressive ability to generate diverse ImageNet $256\times 256$ images across a wide variety of categories. Appendix B.5 show the Failure case of NFIG.
Scaling Up. To validate NFIG’s scaling behavior, we train 310M and 600M parameter models for 55 epochs under computational constraints. The results are shown in Table 3.
Performance of Different Epochs. VAR sets different training epochs for models of different sizes: 200 epochs for 310M, 250 epochs for 600M, 300 epochs for 1B, and 350 epochs for 2B. Limited by computational resources, we focus on training our 310M model. As Table 4 shows, NFIG outperforms VAR at matched epochs and demonstrates superior parameter efficiency. At 200 epochs, NFIG already outperforms VAR-d16 of the same size. With extended training, NFIG-310M achieves performance comparable to or better than VAR-d20-600M (twice the parameters) while using significantly fewer resources.
Table 3: Performance of NFIG with different parameters at 55 epochs.
[…]
NFIG-310M 5.47 224.20 0.7569 0.4914 55 310M 10 1
NFIG-600M 5.07 225.16 0.7184 0.5546 55 600M 10 1.25
@@
Table 4: The performance of NFIG and VAR at different epochs.
[…]
VAR-d16 200 3.55 274.4 0.84 0.51 310M 10 1
VAR-d20 250 2.95 302.6 0.83 0.56 600M 10 1.25
NFIG(ours) 200 3.35 309.2 0.79 0.55 310M 10 1
NFIG(ours) 250 3.16 311.9 0.78 0.56 310M 10 1
NFIG(ours) 300 2.93 325.9 0.79 0.56 310M 10 1
NFIG(ours) 350 2.81 332.4 0.77 0.59 310M 10 1
@@ 4.4 Motivation Verification
Frequency Distribution Analysis. The experimental results in Figure 4 demonstrate the progressive refinement of generated images and the effective capture and synthesis of multi-scale visual features by FR-VAE. The frequency spectrum visualizations reveal the model’s ability to hierarchically incorporate information from low to high frequencies, resulting in generated images with rich details and natural appearance. Appendix B.2 compares the frequency keep ability of NFIG and VAR.
Figure 4: Generated images at different steps 2, 4, 6, 8, 10 of a 10-step process by FR-VAE, with corresponding frequency spectrum. In these spectrograms, brightness (red/yellow) indicates higher frequency energy while darker colors (blue) represent lower energy components. The center of each plot shows low-frequency information, with frequencies increasing radially outward, revealing the evolving distribution during the generation process.
@@ 6 Limitation and Future Work
Our frequency-guided autoregressive image generation approach shows promise, but has limitations. Improving frequency decomposition. A primary issue is the simplistic frequency band division by scale, which inadequately captures information in the first band. Implementing a more rigorous division based on statistical analysis and physical principles would likely enhance NFIG’s performance. Recent advances in beneficial noise theory [43, 44] suggest that properly designed noise can reduce task complexity, which could inform better frequency decomposition and data augmentation strategies [45, 46]. Extension to other modalities. Beyond 2D spatial frequency, future work could extend to video generation by incorporating temporal frequency decomposition, or to 3D object generation where frequency analysis is vital for accurate light field representation. For multi-modal generation, techniques that enhance cross-modal alignment through learnable noise [47] may offer insights for frequency-based fusion strategies. Privacy-preserving generation. Adversarial noise techniques [48] could be integrated with our framework to ensure privacy protection in generated content. Due to computational constraints and time limitations, these promising directions remain unexplored and are left for future investigation.
@@ A.3 Magnitude Spectrum and Power Spectrum
Appendix B: Addition Experiments
VAR sets different training epochs for models of different sizes: 200 epochs for 310M, 250 epochs for 600M, 300 epochs for 1B, and 350 epochs for 2B. Limited by computational resources, we focus on training our 310M model. As the table above shows, NFIG outperforms VAR at matched epochs and demonstrates superior parameter efficiency. At 200 epochs, NFIG already outperforms VAR-d16 of the same size. With extended training, NFIG-310M achieves performance comparable to or better than VAR-d20-600M (twice the parameters) while using significantly fewer resources.
B.1 Details of Loss function
We will provide detailed mathematical formulations of our loss function components and their respective roles in the training process. The total loss function for the FR-VAE (image tokenizer for NFIG) is defined as:
@@ B.1 Details of Loss function
Here, the first two terms represent the reconstruction loss for the image ($I$ vs $\hat{I}$) and and its frequency-guided quantized loss ($\hat{f}$ vs $\hat{f}$), respectively, ensuring fidelity in both pixel and feature. $\mathcal{L}_{p}$ is LPIPS perceptual loss and $\mathcal{L}_{g}$ is gan loss.
For the NFIG Transformer, which predicts the frequency tokens, we utilize a standard cross-entropy loss:
$\mathcal{L}(T,\tilde{T})=-\sum_{i=1}^{n}t_{i}\log(\tilde{t}_{i})$ (11)
This loss is computed between the predicted tokens $\tilde{T}$ and FR-VAE ground truth tokens $T$, ensuring accurate prediction of the quantized frequency representations across all scales.
B.2 Frequency Keep Ability
We are grateful for your encouragement to discuss both successes and challenges. Your suggestion for a frequency analysis was particularly insightful. As you requested, we conducted a frequency-domain comparison between our model (NFIG) and VAR-16.
As requested, we provide frequency-domain comparisons between VAR-16 and NFIG: (1) Power Spectral Density (PSD): Overall frequency fidelity; (2) Frequency Keep Score (FKS): Weighted similarity across High/Mid/Low frequency bands (weights: 0.15, 0.28, 0.57, emphasizing structural low-frequency information).
Our analysis revealed that while both models effectively preserve low-frequency information, and NFIG preserves middle and high frequency information with higher fidelity.
[…]
VAR-16 0.87 79.5% 98.3% 57.6% 48.2%
NFIG(ours) 0.47 87.6% 98.9% 75.3% 66.7%
B.4 Diverse Image Types
To demonstrate broader applicability, we conducted preliminary reconstruction evaluations (FID) of FR-VAE across diverse image types.
Model DTD QRCODE Diagrams Chest-X CelebA-HQ COCO LSUN-Bedroom
FR-VAE 6.86 11.01 21 0.74 3.51 7.51 6.12
B.5 Failure Case
Despite strong overall performance, NFIG occasionally produces visual artifacts. As shown in Figure 6, these include anatomical errors (extra bird leg), texture abnormalities (goldfish patterns), and fine detail loss (bird claws). Red boxes highlight the anomalies. These failures reflect challenges in maintaining semantic consistency across frequency bands. The issues are particularly pronounced for complex structures and fine details. Such limitations are common to frequency-based generation approaches and present opportunities for future improvement.
Figure 6: Failure case for NFIG.